The recovery time in days for 36 patients are given below

8 7 6 4 5 3 7 8 10 7 7

6 4 10 3 6 8 2 5 4 5 3 8

7 4 6 3 7 12 4 3 6 6 9 4

A. Construct a frequency distribution

B. What percentage of recovery days were fewer than 6

C. Find the mean and median number of days to recovery

D. Fin the 20th percentile and the samllr variance


1
Expert's answer
2021-12-14T17:16:30-0500

Least to Greatest Value:

2, 3, 3, 3, 3, 3, 4, 4, 4, 4, 4, 4, 5, 5, 5, 6, 6, 6, 6, 6, 6, 7, 7, 7, 7, 7, 7, 8, 8, 8, 8, 9, 10, 10, 12


A.

value frequency

2 1

3 5

4 6

5 3

6 6

7 6

8 4

9 1

10 2

12 1


B.

"15/35=0.4286=42.86\%"


C.

mean:

"\mu=\sum x_i/n=5.91"


median is the value separating the higher half from the lower half of a data sample

in our case, n = 35, so median is 18th value:

median = 6


D.

sample variance:

"s^2=\frac{\sum (x_i-\mu)^2}{n-1}=5.49"


"P_{20}=(20(n+1)\/100)"th value of the observation

"P_{20}=(20(35+1)\/100)=(7.2)"th value of the observation

= 7th observation + 0.2(8th-7th) = "4+0.2(4-4)=4"